OpenStudy (anonymous):

A rectangular parking lot is 50 ft longer than it is wide. Determine the dimensions of the parking lot if it measures 250 ft diagonally.

OpenStudy (anonymous):

let w=width l=length then l=w+50 Now the diagonal can be found by the pythagorean theorem diagonal=d=250 d^2=l^2+w^2 250^2=(w+50)^2+w^2 Rearranging 0=2w^2+100w-60000 Use quadratic formula and use the positive root as the width must be greater than zero w=150 then l=200

OpenStudy (anonymous):

thank you